# A Grassroots Architecture to Supplant Global Digital Platforms by a Global Digital Democracy

### v1

@@ 1 Introduction
Here we present an architectural alternative to global digital platforms termed grassroots [45, 43]. It is designed to support grassroots platforms that serve the social, economic, civic, and political needs of local digital communities, as well provide for their federation. Grassroots platforms may offer local alternatives to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. The grassroots architecture presented here is a major revision and expansion of an earlier proposal with a similar goal [48].
@@ 2.3 The Blocklace
The blockchain (Figure 2.A, aka block DAG) is a linear data structure in which each block except the first has a hash pointer to its predecessor. It is constructed by agents competing to add the next block to the blockchain. Blockchain consensus protocols such as Nakamoto Consensus [36] resolve the conflicts, or ‘forks’, that arise when two or more agents add their own blocks to the same blockchain, in order to produce the consensual linear blockchain.
@@
The blocklace (Figure 2.B) is a partially-ordered generalisation of the blockchain in which each block has any finite number of signed hash pointers to preceding blocks. It can constructed cooperatively, by agents adding blocks to the blocklace and informing each other of new blocks.

### v2

@@ 1 Introduction
Here we present an architectural alternative to global digital platforms termed grassroots [45, 43]. It is designed to support grassroots platforms that serve the social, economic, civic, and political needs of local digital communities, as well as provide for their federation. Grassroots platforms may offer local alternatives to global digital platforms while operating solely on the smartphones of their members, forsaking any global resources other than the network itself. The grassroots architecture presented here is a major revision and expansion of an earlier proposal with a similar goal [48].
@@ 2.3 The Blocklace
The blockchain (Figure 2.A) is a linear data structure in which each block except the first has a hash pointer to its predecessor. It is constructed by agents competing to add the next block to the blockchain. Blockchain consensus protocols such as Nakamoto Consensus [36] resolve the conflicts, or ‘forks’, that arise when two or more agents add their own blocks to the same blockchain, in order to produce the consensual linear blockchain.
@@
The blocklace (Figure 2.B, aka block DAG) is a partially-ordered generalisation of the blockchain in which each block has any finite number of signed hash pointers to preceding blocks. It can be constructed cooperatively, by agents adding blocks to the blocklace and informing each other of new blocks.